Grade 12th passMechanics

Determine the frequency of revolution of the electron in 2nd Bohr’s orbit in hydrogenatom.

Aditya Pathak
8 Years agoGrade 12th pass

2 Answers

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear Aditya
 
Frequency of revolution = 6.55 * 10^15  z²/n³
Put z = 1 and n = 2..hence
Frequency = 8.187 * 10^14  sec -1
Khimraj
ApprovedApproved Tutor Answer8 Years ago
Frequency of electron revolving in orbit is given by
f = (me4/4n3h3\varepsilon _{0}^{2})(z2/n3)
Now n = 2 and z = 1
f = (6.55*1015)(12/23)
f = 8.187*1014 s-1
Hope it clears.